OpenStudy (anonymous):

how do you put this problem in a equation: General admission tickets to the fair cost $3.50 per person. Ride passes cost an additional $5.50 per person. Parking costs $6 for the family. The total costs for ride passes and parking was $51. How many people in the family attended the fair?

OpenStudy (anonymous):

Help please anyone?

Directrix (directrix):

@rudedog01 Should " The total costs for ride passes and parking was $51. " be " The total costs for admission, ride passes and parking was $51."

Directrix (directrix):

If so, A + R + P = 51 where A is total amount spent on admission, R is total amount spent on rides and P = 6 for parking. 1A + 1R = $9. At $9 for one person and N representing the number of people in the family, 9N + 6 = $51 9N = $ 45 N = 5 people